6. A volume of 310 ml of O2 (g) is collected over water at 20C at a pressure of 738 torr. (a) What is the partial pressure of the O? (b) What would be its volume when dry at STP? (Vapor pressure at 20°C is 17.5 torr) (10) 0 mol N2 exerts a pressure of 800 torr. What is the partial pressure of each gas? (5) 8. A mixture of 2.00 mol H2, 3.00 mol NH3, 64.00 mol Co2, and 5.00 mol N2 was place in a 25.0 L container with a pressure of 750 torr and a temperature of 250 K. What is the total pressure of the container? (10)
Answer #1

6) a) partial pressure of O2 = 738-17.5 = 720.5 torr

    b) Initial volume of O2(V1) = 310 ml

        Initial pressure of O2(P1) = 720.5 torr

        initial temperature (T1) = 293 k

         volume of O2 at STP (v2) = ?

    pressure of O2(P2) = 760 torr

        final temperature (T2) = 273 k


          P1V1/T1 = P2V2/T2

          (720.5*0.31/293) = (760*V2/273)

     v2 = 0.274 l

        = 274 ml

7) TOTAL mol = 2+3+5 = 10 mol

partial pressure of H2(pH2) = xH2*pTotal

                              = (2/10)*800

                              = 160 torr

   partial pressure of N2(pN2) = xN2*pTotal

                              = (5/10)*800

                              = 400 torr

   partial pressure of NH3(pNH3) = xNH3*pTotal

                              = (3/10)*800

                              = 240 torr

8) Total pressure of mixture (P) = nRT/V

       n = total no of mol of mixture = 2+3+64+5 = 74 mol

       R = 0.0821 l.atm.k-1.mol-1

       T = 250 k

       V = 25 l

    P = (74*0.0821*250)/25 = 60.754 atm